Differentiate, y=ln(xlnx)


1
Expert's answer
2021-05-25T16:20:34-0400

"y=ln(xlnx)"


Differentiate w.r.t. x-


"\dfrac{dy}{dx}=\dfrac{1}{xlnx}(lnx+x(\dfrac{1}{x}))"


"=\dfrac{lnx+1}{xlnx}"